Prime Numbers Write a program that will determine whether a set of randomly-generated numbers are prime numbers. To write your program, you will need to use the following programming features: The random library (module) A function A variable that will prompt the user to say how many random numbers the program should generate A for loop with a range function A randint function taken from the random library (module); the randint function takes two arguments – the beginning and the ending of a range An if/else statement print commands, and A calling function. Before you write your program, you might want to outline what needs to be done and write the pseudocode for this program.
Answer #1

Code:

import random
from math import sqrt


def isPrime(num):
    """
    method to determine whether a number is prime or not
    """
    if num < 2:
        return False
    # iterate till sqrt(num) if it is divisible by any other number than it is not prime
    for i in range(2, int(sqrt(num)) + 1):
        if num % i == 0:
            return False

    return True


if __name__ == '__main__':

    # prompt the user to enter how many random numbers he want to generate
    size = int(input("enter how many random numbers you want to generate"))
    # ask user to enter the start and end range in which he want to generate random values
    starting, ending = map(int, input("Enter the starting and ending numbers to generate random numbers").split())
    print(starting, ending)
    numbers = []
    
    # make a list of all the randomly generated numbers 
    for i in range(size):
        numbers.append(random.randint(starting, ending))

    print("The list of randomly generated numbers is : ")
    print(numbers)
    i=0
    # if even one number is not prime  , then display set of generated numbers is not prime
    for i in range(size):
        if isPrime(numbers[i]):
            continue
        else:
            break

    if i < size:
        print("All the numbers in the set of randomly generated numbers are not prime")
    else:
        print("All the numbers in the set of randomly generated numbers are prime")

Screenshots:

pimport random afrom math import sqrt def isPrime(num): if num < 2: return False for i in range(2, int(sqrt(num)) + 1): if nugif main name size = int(input(enter how many random numbers you want to generate)) starting, ending = map(int, input(EnteOutput:

enter how many random numbers you want to generate10 Enter the starting and ending numbers to generate random numbers1 50 1 5